Powdered zinc metal reacts with sulphur (S8) when heated to produce zinc sulphide.
What mass of product can be produced with 25 g of zinc?
What is the actual yield if the percentage yield was 80%?
1
Expert's answer
2016-01-16T08:58:52-0500
8Zn + S8 = 8ZnS
n(Zn) = m/M = 25 / 65 = 0.385 (mol) = n(ZnS)
M(ZnS) = 65 + 32 = 97 (g/mol)
m(ZnS) = n*M = 0.385 * 97 = 37.345 (g)
m’(ZnS) = m(ZnS)*yield = 37.345 * 0.8 = 29.876 (g)

Answer
m’(ZnS) = 29.876 g